# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

### v1

@@ 4.6. Reimagining Social Media’s Potential
Given the nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what social media could be. Therefore, while they acknowledged potential challenges—such as the platform being overly immersive, addictive, or demanding—they largely viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants saw SMH as an inspiring alternative to existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They also appreciated the FI workshops for empowering them to reimagine social media and to be hopeful about its possibilities in ways they had previously thought impossible.
4.6.1. Fostering Hope for the Positive Potentials of Social Media
@@
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “fun” and “novel,” like a “breath of fresh air.”
Beyond features, participants expressed how SMH inspired hope for the future of social media. P10 pointed out that overcoming the design fixation around social media is crucial:
@@
4.6.2. Celebrating Youthful Creativity and Play
SMH allows a playful and engaging social media experience, allowing users to immerse themselves in actively creative activities and self-expression. P06 appreciated this ability to personalize their experience on SMH, stating, “I think… everyone likes sharing parts of themselves.” P21 likened SMH to a “game” due to its creative freedom. They elaborated, “The creative freedom you have over the platform makes it feel more fictional than it is real.”
Participants found the overall experience of SMH to be youthful, fun, and deeply immersive, offering a space that feels both magical and meaningful. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
4.6.3. Enabling Otherwise Inaccessible Experiences
SMH creates opportunities for users to engage in experiences and interactions that may otherwise be out of reach. P09 highlighted how individuals “experiencing stress” or those who may not have access to “really, really beautiful landscape and environments” could walk around and “admir[e]” the “calming” landscape of SMH. P16 envisioned SMH as a space where users could engage in activities like climbing a mountain in Nepal, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
P02 saw SMH as a platform that would “seem more like real life, but better.” They imagined being able to engage with friends in idyllic settings, such as “on the beach all the time.” Furthermore, they believed SMH could strengthen friendships through magical features. In their words, “Since people can’t read others’ emotions in real life, SMH would… be a way for you to connect with your friends on a deeper level.” P18 also appreciated how SMH’s 3D interface facilitated them to interact with “family across the world” in an immersive way. SMH would offer a sense of presence that goes beyond conventional video calls, enabling users to engage with loved ones “not just through a screen.”
@@ 5. Discussion
Figure 3. (a) Current youth perceptions of social media platforms and where they experience meaningful connections, highlighting the disconnect between platforms commonly recognized as “social media” versus where youth actually experience meaningful social connection. The SMH—youth-envisioned ideal social media—is represented by the overlap, combining both aspects. (b) A more nuanced viewpoint clarifies “social media” as an umbrella term for a range of platforms with different primary purposes—friendship-building, entertainment, status, information-seeking, tie maintenance, and more—rather than a few archetypal platforms.
@@
This diagram is divided into two parts, offering insights into how youth perceive social media and how they categorize its purposes. On the left, there are two overlapping circles. The circle on the left represents platforms that youth identify as “social media,” such as Instagram, Facebook, and TikTok. The circle on the right illustrates platforms where youth feel meaningfully connected with friends, including FaceTime, Discord, Minecraft, and Roblox. In the overlapping area falls Social Media at Hogwarts (SMH) youth-envisioned ideal social media that combines the qualities of both categories. On the right, a Venn diagram shows potential categories of social media platforms by their primary purposes, such as “Entertainment-Centered,” “Status-Centered,” and “Information-Seeking-Centered.” “Social Media at Hogwarts” is grouped under “Strong Tie-Centered.”
5.1. “Social Media is 2D”: Youth Perspectives on What Constitutes Social Media
Public discourse increasingly portrays social media as fundamentally detrimental to mental health, friendship, and authentic social interaction. This narrative often equates social media with specific platforms like Instagram, TikTok, or X and the ways in which they happen to be designed today (Kim et al., 2024b).
Our study participants expressed similar perspectives, frequently describing social media as centered around superficial interactions that fail to foster meaningful connections. As illustrated in Figure 3(a), there is a disconnect between platforms that youth commonly identify as “social media” (such as Instagram and TikTok) and the platforms where they actually experience meaningful connections with friends (such as Discord or Minecraft). When envisioning their ideal social platform, participants described something that would bridge this divide—a space combining the deliberate pursuit of social connection they associate with traditional social media with the organic, engaged interactions they naturally experience in gaming and chat platforms.
This misalignment between expectations and experiences around social media is problematic because negative perceptions of social media use can diminish psychological well-being (Lee and Hancock, 2024; Kim et al., 2024a). Youth experience disappointment in seeking deep social connections on platforms where such interactions are scarce. As shown in Figure 3(b), social media serves diverse uses and gratifications—from information seeking and entertainment to status enhancement and network maintenance. Rather than maintaining an entirely negative view of social media confined to certain platforms, it is crucial to understand each platform’s distinct values and limitations.
For instance, critiquing Instagram for not facilitating authentic, deep friendships misses the point; while such connections might occasionally form there, it is not the platform’s central form of engagement. However, it is important to note that platforms aren’t static—their designs evolve, and user engagement patterns shift over time. Instagram, primarily used for personal branding, sometimes facilitates closer communication through private ‘finsta’ accounts. This fluidity in platform use underscores the importance of teaching (Wisniewski et al., 2012a; Wisniewski, 2018a; The Learning Network, 2025) youth to be more discerning about the connections they seek and what’s possible on each platform rather than enforcing anxiety-inducing narratives about social media’s harmful effects.
@@ 5.2. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings confirm prior studies that while the youth are largely dissatisfied with social media and how they spend their time on archetypal social media platforms, they do not necessarily want to reduce their social media usage—rather, they seek platforms that facilitate deeper, more meaningful connections. Our participants identified four key elements that could transform social media from a space of passive consumption to one of genuine connection:
@@
    Presence and Immersion. Youth envision platforms that foster genuine emotional connection through real-time presence. While current platforms feel “flat” and emotionally disconnected, participants valued synchronous interactions that create a sense of co-existence. Even simple activities like sharing quick updates about eating ice cream or walking outside, or simply existing in a shared virtual space were seen as meaningful when done with immersive co-presence.
@@
    Supporting Individual Preferences. Youth desire more natural and diverse ways to express themselves. Participants envisioned platforms where they could show their personality through both active sharing and subtle signals, similar to how personality naturally emerges in physical spaces. From memory-sharing to customizing their virtual spaces and avatars, they want to express themselves spontaneously and casually rather than feeling constrained to carefully curated posts.
@@
    Intuitive Social Navigation. Meaningful social media experiences require design that mirrors how relationships naturally unfold. Youth value platforms that enable organic relationship growth through proximity and repeated low-intensity interactions, similar to how sitting in the same classroom gradually builds connections. They also seek designs that replicate the subtle, embodied norms of physical spaces—where spatial arrangements naturally signal degrees of privacy and social contexts, much like how a bedroom implies different privacy expectations than a living room.
@@
    Playful Engagement. Youth emphasize that meaningful connection often grows through joy and shared experiences, yet they find current platforms overly serious and high-stakes. They envision platforms that encourage whimsical exploration and casual interaction—where they can discover hidden objects, explore virtual landscapes, or simply hang out in a “third place.” This playfulness creates low-stakes opportunities for friendship development beyond the performative nature of mainstream social media.
@@
Our participants envision environments where relationships deepen naturally through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement—challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
@@ 5.3. Spatial Integrity for Meaningful Social Connections in Online Environments
In the design of virtual environments that foster meaningful social connections, spatial cognition—the way we perceive and interpret space—remains critically relevant. Even in virtual spaces, human interactions are shaped by spatial affordances, mirroring many aspects of physical-world interactions. While virtual spaces need not replicate physical environments exactly—indeed, they can transcend physical limitations in some ways while falling short in others—certain spatial principles remain fundamental to how humans build and maintain relationships. This relevance is reflected in our pervasive use of spatial metaphors when describing virtual interactions, such as “chat rooms,” “cyberspace,” or “folders.” While some platforms like Discord and social VR applications successfully implement spatial design elements, mainstream social media platforms often lack these affordances. Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces.
@@
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offers only transient content streams. Our participants emphasized the importance of dedicated areas—like communal lounges or study spaces—where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
@@
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity—a user’s feed” is meant to be personal space yet lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
@@
These four dimensions of spatial integrity—presence, composition, configuration, and depth—are critical for designing virtual environments that support meaningful social connections. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram that we explored in Section 5.2. As spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).)
The concept of spatial integrity offers a framework for understanding why many social media platforms feel disconnected and unsatisfying. By thoughtfully incorporating these principles, we can design online environments that better align with human social-spatial cognition, addressing the limitations our participants identified and better supporting meaningful social interactions and the organic development of relationships.
@@ 5.5. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, potentially biasing them toward favorable attitudes toward incorporating spatial elements or online environments that center play. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research could empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him—including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.

### v2

@@ 4.7. Expanding Horizons
Given the speculative nature of the FI co-design workshop, SMH embodies the youth’s aspirations and values for what interactions on social media platforms could embody. Therefore, while they acknowledged potential challenges, such as the platform being over-immersive or distracting, they overwhelmingly viewed these downsides as worthwhile trade-offs for its many benefits. Overall, participants described SMH as an inspiring alternative or expansion of existing social media, offering opportunities for self-expression and the ability to connect in ways that felt deeply personal and intentional. They saw SMH not just as a better platform, but as an inspiring redefinition of social media’s purpose and potential.
4.7.1. Breaking Free from Platform Monotony with Novel Interactions
@@
Participants also appreciated how SMH offered a break from the monotony of existing platforms. P11 pointed out that most social media platforms share a common “standard” with repetitive features like “feed, messages, settings,… profile.” In contrast, SMH introduced new types of interactions that deviated from familiar patterns. P02 described SMH as “a combination of Discord and BeReal” that brings together casual sharing and topic-based discussions. Similarly, P17 envisioned their SMH as “more interactive, kind of like if Discord and Instagram were combined,” where they could enjoy the interactions of Discord along with profiles and the ability to post content like on Instagram. P22 highlighted that on other platforms, there is limited opportunity for customization or decoration, explaining, “I just have a profile picture.” However, on SMH, they could “really be creative” and “make something that [they] haven’t made anywhere else,” which made the experience feel “novel,” like a “breath of fresh air.”
For some participants, SMH’s novelty lay in giving themselves permission to prioritize enjoyment over seriousness. P14 described initially defaulting to the idea of a more “mature” 2D version, but later embracing the playful, immersive 3D version as more personally meaningful and fun. For example, P14 initially leaned toward preferring a 2D version of the platform but shifted their preference to the 3D version, reasoning:
4.7.2. Inspiring Hope and Expanding the Scope of Social Media Interactions
Through the FI process, participants realized they were not limited by the structures of existing platforms. Instead, they felt empowered to imagine social technologies that center joy, creativity, and authentic human connection, qualities they had once believed were beyond the reach of digital spaces. P10 pointed out that overcoming the design fixation around social media is crucial:
@@ 5. Discussion
Figure 2. 3D Gamer Youth’s Vision of Ideal Social Media Interactions
@@
A Venn diagram comparing two categories of online platforms. The left circle represents platforms focused on social networking and broadcasting, with icons for Instagram, Facebook, and TikTok. The right circle represents platforms where youth engage in meaningful connections, with icons for Zoom, Minecraft, Discord, and Roblox. In the overlapping center, labeled “Social Media at Hogwarts (SMH)”, the diagram highlights youth’s vision of an ideal platform that combines the social visibility of networking platforms with the depth of meaningful connection found in gaming and community spaces.
Figure 2 captures how participants described their vision for ideal social media as something that pulls benefits from both types of platforms they already use. On one side are social networking and broadcasting platforms like Instagram, Facebook, and TikTok, which make it easy to stay socially visible but often feel shallow or performative. On the other side are platforms like FaceTime, Minecraft, Discord, and Roblox, where youth experience more organic and engaging connections, but where social interaction is often secondary to other goals like gaming or productivity. Our participants imagined “Social Media at Hogwarts (SMH)” as a platform that blends these strengths, creating a space where meaningful social connection effectively surfaces not just a peripheral effect, but a key focus of the platform.
@@ 5.1. Youth Seek More Meaningful Connections, Not Less Social Media Minutes
Our findings build on prior research showing that while youth are often dissatisfied with archetypal social media platforms, their goal is not to reduce time spent connecting online. Instead, they seek spaces that support deeper, more intentional social interactions. Participants described social connections that feel genuine, fulfilling, and emotionally present and not performative. The changes in social media platform design that they sought were as follows:
@@
    Immersive Experiences Through Spatial Elements. Youth expressed a strong desire to move beyond “flat” social media toward spatially immersive experiences. They imagined platforms where relationships could unfold through presence in shared 3D environments, not just posts on a feed. They described the value of feeling physically co-present with others, such as walking together in a virtual world or talking in a space that feels alive and embodied. They also envisioned personal spaces and community-centered landscapes, such as personal homes or neighborhoods that mirror real-world social structures. Several imagined futuristic forms of presence, such as those of holographic displays or VR interfaces that make interacting feel natural and embodied, reducing the sense of separation created by screens.
@@
    Organic and Intentional Social Interactions. Participants described meaningful connections as unfolding through casual, low-pressure, and community-oriented interactions. They valued flexible and safe spaces for direct conversations, where people could gather in protected and themed rooms to get to know each other more personally. They also emphasized the importance of community gathering places and shared activity spaces, such as areas to play games, collaborate on projects, or simply hang out. Visions of these spaces encompassed both low-intensity ambient co-presence, such as sitting together quietly while listening to music, and public third places that enable serendipitous encounters with new people. Participants also described the value of effortful and intentional navigation, preferring to move through spaces purposefully rather than having content or people pushed to them by algorithms.
@@
    Expressive and Lower-Stakes Sharing. Participants sought ways to express their identities naturally and casually, without the pressure of curating formal posts. They valued personal area and avatar customization as ways to reflect their personality through how they decorate spaces or design their avatars. They also envisioned ways to share real-time states and updates, such as showing what they are doing in the moment, fostering a sense of immediacy and connection. Participants described the importance of sharing multisensory or emotional memories, imagining ways to let others experience meaningful moments, like a special memory or sensory detail, in richer and more embodied ways than existing media allows.
@@
    Granular, Intuitive, and Playful Privacy Mechanisms. Participants emphasized the need for privacy that feels natural and understandable, not overwhelming or rigid. They described space-based privacy, where the type of space, such as a public plaza or a personal bedroom, intuitively signals boundaries. They also described playful privacy mechanisms, such as secret knocks or hidden rooms, that make managing access feel engaging and enjoyable. Participants expressed the importance of contextual privacy, with the ability to adjust what different people see in different contexts or moments. They wanted to feel fully in control of their visibility, including the ability to become invisible or undiscoverable to people they block or want to avoid.
  5. (5)
    Refocusing Interaction Priorities for Meaningful Engagement. Participants expressed frustration with how mainstream platforms prioritize endless content streams, celebrity updates, and commercial content over personal connections with people they care about. They described the value of intentional content consumption, where they could choose what they wanted to engage with rather than being fed endless streams. They also advocated for reducing celebrity and commercial influence, preferring to focus on posts from close connections. Some participants described using spatial representations of relationships, such as placing their closest friends’ spaces physically nearer to them in the platform’s landscape. Across these ideas, participants emphasized the value of purposeful, meaningful use of time, where interaction feels worthwhile because it supports relationships, not just engagement metrics or passive scrolling.
@@
Our participants envision environments where relationships deepen organically through shared experiences and deliberate engagement rather than through metrics-driven interactions. This vision suggests that platforms prioritizing meaningful connections could redefine online social interaction while maintaining user engagement, challenging the assumption that profitable social media must rely on superficial, compulsive engagement strategies.
@@ 5.2. Spatial Integrity for Meaningful Social Connections in Online Environments
Concept Description Supported Interactions
Spatial Presence The ability to perceive and navigate a virtual space, including awareness of other movement, and presence. Gaming and VR platforms excel at this, whereas mainstream social media provides limited spatial affordances for awareness and intentional navigation. Immersion and embodiment (Section 4.1); memory and experience sharing (Section 4.4.3); co-existing in surreal landscapes (Section 4.6.3)
Spatial Configuration The presence of fixed/defined spatial structures that facilitate interaction. Platforms like Discord create stable communal spaces with clear channels and streams. Users value this for fostering deeper relationships. Neighborhood structures (Section 4.3.1); prioritizing closer connections (Section 4.3.3); creating quality conversational spaces (Section 4.3.4); imposing clearer boundaries (Section 4.3.5); deliberate navigation and content discovery (Section 4.5.1)
Spatial Composition The segmentation of a social network into defined areas and pathways. Clear boundaries help manage social norms and interactions, with current platforms blurring personal and public spaces. Building comfort through ambient presence (Section 4.2.2); surfacing third places (Section 4.3.2)
Spatial Depth The amount of layered, rich information and space supporting richer, more intuitive posting and commenting. 3D social media limits expression to explicit actions like posting and commenting. Leveraging spatial cues for intuitive privacy control (Section 4.1.2); indirect self-expression through customization and artifact display (Section 4.4.1); memory and experience sharing (Section 4.4.3)
Table 2. Dimensions of Spatial Integrity in sociotechnical systems in supporting meaningful social connections.
Through our interviews, we identify four key dimensions of what we term spatial integrity to support meaningful social interactions in online spaces:
@@
    Spatial Composition: The Meaning of the Space. Spatial composition refers to “the presence of fixed places… that make interaction possible or likely” (Small and Adler, 2019). Platforms like Discord successfully create persistent “third places” through dedicated servers and channels that serve as anchors for social connection (JaeWon et al., 2025), while many archetypal social media typically offer only transient content streams. Our participants emphasized the importance of dedicated areas like communal lounges or study spaces where friends could regularly gather, noting how such persistent spaces enable relationships to develop depth through shared memory and history.
@@
    Spatial Configuration: How the Space is Organized. Spatial configuration is “the segmentation of space into subunits with physical boundaries and pathways between them” (Small and Adler, 2019). Clear boundaries and pathways help people recognize territories and manage social norms. Current platforms lack such clarity; a user’s “feed” is meant to be a personal space yet often lacks clear boundaries, while profiles can be accessed without users’ awareness. This absence of defined pathways for close friend interactions and boundaries for access control makes social media spaces feel disorderly or invasive, undermining meaningful social boundaries.
@@
According to our participants, these dimensions serve as the foundation for designing online spaces that effectively support the kind of social connections they desire. The relationship between these dimensions and the interactions they support is laid out in Table 2.
These four dimensions of spatial integrity—presence, composition, configuration, and depth—provide a framework for designing online spaces that support meaningful social connections. By prioritizing these dimensions, platforms can create environments where the kind of meaningful interactions our participants desire can flourish. While some platforms have successfully incorporated these elements, mainstream social media often lacks these affordances, contributing to the dissatisfaction youth feel toward platforms like Instagram. Especially as spatial computing technologies gain prominence, there is an opportunity to integrate spatial integrity into the design of future platforms. (It is important to note that while spatial depth is generally more pronounced in 3D environments, the other dimensions of spatial integrity can be effectively achieved in 2D spaces, as demonstrated by platforms like Discord (JaeWon et al., 2025).) By thoughtfully incorporating these principles, platforms can create spaces that foster genuine connection and community, aligning with the desires expressed by our participants.
@@ 5.3. Overcoming Cognitive Fixation and Doom Narratives Through Fictional Inquiry
There are societal concerns about social media’s impact on youth well-being (Oth, 2023). Rather than addressing these concerns constructively, public discourse has increasingly framed social media as a societal evil (Wike et al., 2022), particularly in the U.S., where it is often blamed for issues ranging from democratic instability to what some call the “Cause of the Mental Illness Epidemic” (Haidt, 2023). This narrative has prompted restrictive policy responses worldwide, such as the “Stop Addictive Feeds Exploitation (SAFE) For Kids Act.”
However, taking social media away from young people or framing it as inherently harmful oversimplifies a much more complex reality. Despite public concern about social media’s role in youth mental health, the causal evidence remains inconclusive, and youth themselves report a mix of positive and negative experiences (Vogels, 2023). At the same time, research shows that fear-based policies and alarmist narratives can do harm of their own, amplifying anxiety, mistrust, and feelings of helplessness among youth, rather than supporting constructive engagement (Kim et al., 2024a; Lee and Hancock, 2024). This misalignment suggests that the solution lies not in rejecting social media altogether, but in asking what these platforms could become if they were designed to better reflect the kinds of social connection that youth actually seek. Rather than accepting the limitations of existing platforms as inevitable, there is an opportunity to rethink their underlying priorities and architectures.
@@ 5.4. Limitation and Future Work
Our study has several key limitations to consider. First, all participants had prior experience with 3D gaming platforms, so the findings do not inherently generalize to youth who are not familiar with such environments. Second, although we were cognizant of how fixed perceptions around “social media” might constrain ideation, our continued use of this terminology throughout the interviews may have inadvertently reinforced existing mental models rather than encouraging completely novel conceptualizations. Third, while the Harry Potter narrative framework successfully transported participants into a magical design space, our specific scenario about a Muggle-born student introducing social technology sometimes led participants to focus on wizard-Muggle divisions rather than the intended focus on connection across distance. This suggests opportunities to refine the narrative approach.
Future research should empirically test the spatial integrity framework to validate whether incorporating its different dimensions indeed fosters more meaningful social connections in online spaces. This could involve prototyping and evaluating platforms that systematically vary presence, composition, configuration, and depth.
Another area for future work stems from participants’ surprising openness toward the idea of Dobby, a house elf, as a personal assistant. They imagined Dobby having limitless magical powers to help maintain their personal spaces and said they would share almost everything with him, including personal details they would not share with friends. Even when interviewers mentioned that Dobby might report back to the house elf society (a metaphor for social media companies), participants did not seem concerned about privacy and remained willing to share. This is particularly striking given that participants were otherwise skeptical of AI. This raises interesting questions about how youth might engage with AI when it is framed within a familiar and relatable narrative. It suggests that embedding AI in natural, empathetic contexts might lead people to trust and interact with it far more than they otherwise would. Future research should delve into the privacy implications of this dynamic and explore how narrative framing could influence users’ relationships with AI in ways that are not immediately obvious.